OpenStudy (kamille):

Hello, I need someone to check it for me, if I have done everything correct? \[5\cos ^{2}a-7cosa-6=0\] Let's say that cosa=m \[5m ^{2}-7m-6=\] \[D=(-7)^{2}-4*5*(-6)=49+120=169\] \[m _{1}=\frac{ 7+13 }{ 10 }=2\] \[m _{2}=\frac{ 7-13 }{ 10 }=-0.6\] \[cosa=2 \] it is not the sollution and \[cosa=-0.6 \] So, the problem starts here. Is -0.6 an answer or do I need to write like this cosa=-0.6 a=\[\alpha=\pm \arccos \alpha +2 \pi k\]

OpenStudy (kamille):

@stamp

OpenStudy (sirm3d):

i agree with your answer.

OpenStudy (kamille):

well, is -0,6 an answer?

OpenStudy (anonymous):

you would have to sub back in

OpenStudy (anonymous):

so where you got \[m_1,m_2\] would be \[cos(\alpha_1),cos(\alpha_2)\]

OpenStudy (anonymous):

then find what alpha is by taking the arcs of each

hartnn (hartnn):

is the Question , give general values of 'a' , or just solve for 'a' or any interval is given ? also, wouldn't it be \(a= \arccos(-0.6)\pm 2\pi k\)

OpenStudy (kamille):

No, there is no interval: just solve the equation

OpenStudy (anonymous):

it'd be a general solution of the values

OpenStudy (sirm3d):

\[\alpha=\pm arccos(-0.6) + 2k\pi\]

hartnn (hartnn):

so, you have to write an general expression, \(a= \arccos(-0.6)\pm 2\pi k\)

OpenStudy (kamille):

and the most confusing part of the problem is the second part: "Find sin alpha, if we know that alpha belongs to (pi; 3pi/2). I know i can solve sin aplha from sin^2a+cos^2a=1, but it would be much more easier if I just leave a=-0.6 in the first part.

hartnn (hartnn):

why \(\pm\) is with arccos and not with 2kpi ?

OpenStudy (sirm3d):

there are two sets of solutions, one \[\arccos (-0.6)+2k\pi\] and the other \[-\arccos(-0.6)+2k\pi\]

OpenStudy (sirm3d):

@Kamille 's factorization is correct, @stamp

hartnn (hartnn):

but a=-0.6 is not correct solution for 1st part.

OpenStudy (kamille):

so, do I need to find value of alpha in interval of {pi;3pi/2)?

hartnn (hartnn):

(pi; 3pi/2) means 3rd quadrant, so when you find sin a from that identity, just add a negative sign because sin is negative in 3rd quadrant.

OpenStudy (stamp):

\[5x^2-7x-6\]\[(5x+3)(x-2)\] \[5cosx+3=0\]\[x=cos^{-1}(-3/5)\] \[cosx-2=0\]\[x=cos^{-1}(2)\] How about that

OpenStudy (sirm3d):

the range of cosine is between -1 and +1, @stamp

OpenStudy (sirm3d):



OpenStudy (sirm3d):

that's why there are two values of alpha corresponding to arccosine -0.6. \[a=\pm \arccos(-0.6) + 2k\pi\quad,\text{k an integer}\]

OpenStudy (kamille):

So i have: a=\[a=\pm \arccos (-0.6) + 2 \pi k\] Can someone remind me how to get radians of 0.6?

OpenStudy (sirm3d):

you have to use your calculator.

OpenStudy (kamille):

well, at school we had a formula (which I have forgotten) like if you have 1/2 you can divide(or multiply) something by pi etc, do you know it?

OpenStudy (sirm3d):

i think the formula works only for special angles.

hartnn (hartnn):

could you do 2nd part, finding sin a ? and numerical value of arccos(-0.6) ?